# A functional shunt in the umbilical cord: the role of coiling in solute and heat transfer

### v1

@@ I Introduction
Figure 1: The vascular structure of the umbilical cord. (a) A schematic of the helical structure of the umbilical cord that connects the fetus to the placenta; the pitch $2\pi/\Omega$ represents the average length of a coil relative to the radius of the cord; $\Omega$ is the dimensionless helicity parameter (proportional to the UCI). (b) A histological image of the cross-section of a healthy human cord (adapted from [7] under CC BY-NC 4.0); the vascular cord configuration is characterised by the angle between the umbilical vein and the first artery ($\theta_{1}$, measured relative to the $x$-axis that passes through the centre of the cord and the centre of the vein), the angle between the two arteries $\Delta\theta$, and the artery–vein separation distances $d_{1},\,d_{2}$ normalised by the cord radius. (c) Schematic illustrations of the hypercoiled and hypocoiled umbilical cords.
@@
The objective of this study is therefore to quantify the diffusive coupling of umbilical vessels in the human cord for different solutes and heat by combining a theoretical model with anatomical data. Our model moves beyond the coiling index by encapsulating the complexity of three-dimensional helical vascular configurations in a computationally efficient and interpretable manner, allowing us to define key geometric determinants of functional shunting in the cord for a wide range of solutes. We further contextualise our results for specific structural features of human umbilical cords, based on ex vivo histology and in vivo ultrasound imaging, and show how typically observed vascular configurations tend to minimise the heat exchange between the cord’s vessels. Our model also predicts that cord helicity can enhance oxygen availability to the cord tissue surrounding the umbilical vessels, particularly in the presence of tissue metabolism. The developed methodology provides a building block for future fetal–placental–maternal models, enabling more direct hypothesis testing of structure–function relationships in obstetric pathophysiology, as well as in other complex exchange organs.
@@ II.1 Modelling solute exchange between helical vessels
We model the umbilical cord as a straight circular cylinder that hosts three helical vessels (one vein and two arteries). Each vessel has a circular cross-section in the plane orthogonal to the axis $z$ of the cylinder and a constant pitch $2\pi/\Omega$, defined as the length of one complete helical turn parallel to the $z$-axis of the cord, measured relative to the cord’s radius (see Fig. 1a). The dimensionless helicity parameter $\Omega$ is proportional to the UCI. The vascular configuration in the cross-section of a cord (Fig. 1b) can be characterised by the relative distance $d$ between the umbilical arteries and the vein (scaled by the cord’s radius) and the angles between the vein and each of the two arteries $\theta_{1}$ and $\theta_{2}$ (measured with respect to an axis connecting the centre of the cord to the centre of the vein). For simplicity, we assume that both umbilical arteries are of the same size and are the same distance $d$ from the vein (so that $d_{1}=d_{2}=d$ in Fig. 1b).
@@
We model the transfer of solutes (such as oxygen and heat) in the cord tissue ${\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}$ by a steady diffusion-uptake process and assume the cross-sectional solute concentration to be approximately uniform in each umbilical vessel. We neglect possible differences in solute concentration between the umbilical arterial blood and the amniotic fluid. The governing equations (see Supplement S1 for more details) in normalised dimensionless variables describing solute transport in the umbilical tissue are given by
@@
$\nabla^{2}C$ $=\alpha\left(C+1/\beta\right)\quad\text{in }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}\,,$ (1a)
$\eta\,C+(\mathbf{\hat{n}}\cdot\nabla C)$ $=0\quad\text{on }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{c}}\,,$ (1b)
$\mathbf{\hat{n}}\cdot\nabla C$ $=0\quad\text{on }\;{\color[rgb]{0,0,0}\mathcal{I}}_{\text{f,\,p}}\,,$ (1c)
@@
[…]
We reduce the three-dimensional problem (1) to two dimensions by exploiting the helical coordinate transformation $X=x\cos{\left(\Omega z\right)}+y\sin{\left(\Omega z\right)}$, $Y=-x\sin{\left(\Omega z\right)}+y\cos{\left(\Omega z\right)}$, $Z=z$, where $z$ measures the distance along the cord’s axis (see Supplement S1 for more details). This transformation allows us to study the three-dimensional effects of helicity in the cross-sectional plane of the cord by solving for $C(X,Y)$, the field that is invariant along the length of the cord relative to a rotating coordinate system. The dimensionless solute flux $N$ per unit length of the cord exchanged between the UV and the two UAs, in the absence of tissue metabolism ($\alpha=0$), and the solute flux per unit length taken up by cord tissue are then given, respectively, by
$N{\color[rgb]{0,0,0}(\Omega,\alpha,\beta,\eta)}=\int_{{\color[rgb]{0,0,0}\partial\mathcal{I}}_{\text{v}}{\color[rgb]{0,0,0}\,\cap\,\mathcal{P}}}\mathbf{\hat{n}}\cdot\left(\nabla_{\perp}C+\Omega^{2}\mathbf{H}\right)\mathop{}\mathrm{d}s\quad\;\text{and}\;\quad N_{\text{u}}=\int_{{\color[rgb]{0,0,0}\mathcal{I}}_{\text{t}}{\color[rgb]{0,0,0}\,\cap\,\mathcal{P}}}\!\left(C+1/\beta\right)\!\mathop{}\mathrm{d}A\,,$ (2)
[…]
$\mathrm{Da}=\frac{D_{\text{t}}L\,N(\Omega,\alpha,\beta,\eta)}{B\,Q}\,,$ (3)
where $D_{\text{t}}$ is the solute diffusivity in tissue, $L$ is the cord length, $Q$ is the total umbilical flow rate and $B$ is an advection-facilitation parameter that accounts for haemoglobin binding of certain solutes, such as oxygen; see Supplement S8 for more details and estimates of $\mathrm{Da}$ for oxygen and heat exchange.
As the metabolic rate $\alpha$ increases, the region surrounding the vessels’ lumens exhibits thin boundary layers and steep concentration gradients. We exclude the effects of these highly localised boundary layers by calculating the extravascular uptake flux within the cord tissue outside a thin strip of thickness $h$ around each vessel (based on typical values for the thickness of the umbilical vessel walls [20]; see Tables S1, S2 in the Supplement). The solute distribution and the corresponding integral fluxes (2) are computed via a finite-element solver of COMSOL ${}^{\text{\texttrademark}}$ 6.0. We validated the use of the 2D field $C(X,Y)$ against direct computationally-expensive solutions of the 3D problem (1) (see Supplement S5 for more details).
@@ II.2 Image analysis of histology and ultrasound data
Histology and ultrasound images of human cord’s cross-sections were taken from published literature [7, 21, 22, 23, 18, 24, 25, 26]. The histology data (stained with hematoxylin and eosin; see Fig. 1b) corresponds to normal pregnancies, with gestational age varying from 37 to 42 weeks; ultrasound data included cords from the third trimester (see Supplementary Table S4 for more details on data sources).
@@ III.1 Cord coiling and vascular proximity promote solute and heat shunting
Figure 2: The impact of cord helicity and vascular proximity on solute exchange. (a) Relative exchange flux $N_{\mathrm{rel}}$ per unit length (4) of the umbilical vein vs. cord helicity $\Omega$ for different UA-UV separation distances $d$, with $\alpha=0,\,\eta=0$. For weakly coiled cords, the solute (heat) exchange amplification scales as $\Omega^{2}$ (relative to the uncoiled vessels of the same length). Concentration fields for four different vessel configurations and three-dimensional diagrams show a hypocoiled ($\Omega$=0.75) and hypercoiled ($\Omega$=3.77) cord, at parameters highlighted with squares on the graph. The colour bar shows solute concentration in normalised units (venous concentration is unity and arterial concentration is zero). (b) Computed exchange flux per unit length of the vein $N$ for small separation $d$. In the limit of small $d$, the leading-order flux is $\mathcal{O}(d^{-1/2})$. Here $\eta=0$. (c) Exchange flux $N$ and relative exchange flux $N_{\mathrm{rel}}$ for different values of the exchange parameter $\eta$ in the cord boundary condition (1b) for $d=0.33$. Other geometric parameters in (a-c) were fixed: $R_{v}=0.25$, $r_{v}=0.25$, $r_{a}=0.12$, $\theta_{1}=0.8\pi$, $\theta_{2}=1.2\pi$.
@@
$N_{rel}=\frac{N(\Omega{\color[rgb]{0,0,0},\alpha,\beta,\eta})-N(0{\color[rgb]{0,0,0},\alpha,\beta,\eta})}{N(0{\color[rgb]{0,0,0},\alpha,\beta,\eta})},$ (4)
assuming initially that $\alpha=\eta=0$. Small $\Omega$ corresponds to weakly coiled cords, while a large $\Omega$ corresponds to a tightly coiled cord. This is shown in the 3D representations of one complete coil for $\Omega=0.75$ and $\Omega=3.77$ (Fig. 2a, insets). For large $\Omega$, $N_{rel}$ exceeds unity, implying that the flux can be twice as large as the flux for straight vessels, because tightly coiled cords provide a larger internal surface area for exchange to take place. The concentration field for $\Omega=3.77$, $d=0.3$ (Fig. 2a, inset, top right) varies weakly around the cord perimeter, showing how coiling promotes direct solute exchange between vessels. In contrast, the flux for weakly coiled vessels is similar to the flux for straight vessels. Fig. 2(a) shows that the first correction to the straight vessel flux is of order $\Omega^{2}$. A weak helicity $\Omega\ll 1$ approximation for flux can therefore be obtained by perturbing the three-dimensional problem (1a) in powers of $\Omega^{{\color[rgb]{0,0,0}2}}$.
Since in the problem for $C(X,Y)$ the helicity term appears as $\Omega^{2}$, there is no evidence of chirality in the concentration fields in Fig. 2(a). In the Supplement (Fig. S7b,c), we confirm that two configurations with opposite chirality have the same cross-sectional concentration fields. The position of the arteries in the configuration considered in Fig. 2(a) is symmetrical relative to the $x$-axis. This leads to the symmetry $C(X,Y)=C(X,-Y)$. The Supplement (Fig. S7a) shows concentration fields with increasing helicity for asymmetric configurations. Breaking the symmetry introduces distortion in the concentration fields, but the trend in relative flux is the same as in Fig. 2(a).
Another geometric parameter that contributes to the exchange of solutes is the separation distance $d$ between vessels. The flux is larger for the umbilical vessels that are closer together because concentration gradients are steeper, as shown in Fig. 2(a), colour map inserts. Indeed, Fig. 2(b) illustrates how, as $d$ gets smaller, the flux in the gap between the vessels of $\mathcal{O}(d^{-1/2})$ dominates over the flux outside the gap (see Supplement S3 for more details), and the weak helicity correction of $\mathcal{O}(d{\color[rgb]{0,0,0}{}^{1/2}}\,\Omega^{2})$ provides a good approximation ($d=0.04$; Fig. 2a). On the other hand, for larger $d$, the effect of helicity becomes more dominant, as demonstrated by the larger excess flux $N_{rel}$ (4) relative to the flux in the uncoiled geometry.

### v3

@@ Abstract
umbilical cord, vascular configuration, helicity, oxygenation, diffusive solute transport, heat exchange
@@ I Introduction
Figure 1: The vascular structure of the umbilical cord. (a) A schematic of the helical structure of the umbilical cord that connects the fetus to the placenta; the pitch $2\pi/\Omega$ represents the average length of a coil relative to the radius of the cord; $\Omega$ is the dimensionless helicity parameter (proportional to the UCI). (b) A histological image of the cross-section of a healthy human cord (adapted from [7] under $C$ BY-NC 4.0); the vascular cord configuration is characterised by the angle between the umbilical vein and the first artery ($\theta_{1}$, measured relative to the $x$-axis that passes through the centre of the cord and the centre of the vein), the angle between the two arteries $\Delta\theta$, and the artery–vein separation distances $d_{1},\,d_{2}$ normalised by the cord radius. (c) Schematic illustrations of hypercoiled and hypocoiled umbilical cords.
@@
The objective of this study is therefore to quantify the diffusive coupling of umbilical vessels in the human cord for different solutes and heat by combining a theoretical model with anatomical data. Our model moves beyond the coiling index by encapsulating the complexity of three-dimensional helical vascular configurations in a computationally efficient and interpretable manner, allowing us to define key geometric determinants of functional shunting in the cord for a wide range of solutes. We further contextualise our results for specific structural features of human umbilical cords, based on ex vivo histology and in vivo ultrasound imaging, and show how typically observed vascular configurations tend to minimise heat exchange between the cord’s vessels. Our model also predicts that cord helicity can enhance oxygen availability to the cord tissue surrounding the umbilical vessels, particularly in the presence of tissue metabolism. The developed methodology provides a building block for future fetal–placental–maternal models, enabling more direct hypothesis testing of structure–function relationships in obstetric pathophysiology, as well as in other complex exchange organs.
@@ II.1 Modelling solute exchange between helical vessels
Parameter Notation Value (cm) Source
Arterial radius $r^{*}_{a}$ 0.060.20.06-0.2 [20]
Venous radius $r^{*}_{v}$ 0.10.40.1-0.4 [20]
Cord radius $R^{*}_{c}$ $0.2-1\,^{\dagger}$ [21]
Venous helical radius $R^{*}_{v}$ 00.80-0.8 [14]
Venous wall thickness $h^{*}_{v}$ 0.030.050.03-0.05 [22]
Arterial wall thickness $h^{*}_{a}$ 0.050.080.05-0.08 [22]
Cord pitch $2\pi/\Omega^{*}$ 3143-14 [23]
Cord length $L^{*}$ 5010050-100 [18]
Artery–vein distance $d^{*}$ 0.040.40.04-0.4 Estimated
Table 1: Key geometric parameters of the model. Stars denote dimensional quantities. The pitch was calculated as the inverse umbilical coiling index (UCI) [23], and the artery–vein distance was estimated as described in Supplement S6. All the dimensionless lengths in the model are scaled on $R^{*}_{c}=1$ cm.
@@
We model the umbilical cord as a straight circular cylinder that hosts three helical vessels (one vein and two arteries). Each vessel has a circular cross-section in the plane orthogonal to the axis $z$ of the cylinder and a constant (and common) pitch $2\pi/\Omega$, defined as the length of one complete helical turn parallel to the $z$-axis of the cord, measured relative to the cord’s radius (see Fig. 1a). The dimensionless helicity parameter $\Omega$ is proportional to the UCI. The vascular configuration in the cross-section of a cord (Fig. 1b) can be characterised by the relative distance $d$ between the umbilical arteries and the vein (scaled by the cord’s radius) and the angles between the vein and each of the two arteries $\theta_{1}$ and $\theta_{2}$ (measured with respect to an axis connecting the centre of the cord to the centre of the vein). For simplicity, we assume that both umbilical arteries are of the same size and are the same distance $d$ from the vein (so that $d_{1}=d_{2}=d$ in Fig. 1b).
@@
We model the transfer of solutes (such as oxygen) and heat in the cord tissue $\mathcal{I}_{\text{t}}$ by a steady diffusion-uptake process and assume the cross-sectional solute concentration to be approximately uniform in each umbilical vessel. While blood flow in the umbilical vein is approximately steady, with weak fluctuations [24], arterial flow is pulsatile [12]. However, the diffusion (with tissue diffusivity $D^{*}_{t}$) of oxygen and heat between the vessels is much slower ($(d^{*})^{2}/D^{*}_{t}\sim 10-10^{3}$ s, for $d^{*}\sim 1$ mm; see Table 1 and Supplement) compared to the typical fetal heart-rate timescale ($10^{-1}-1$ s), justifying the steady-transport approximation. We also neglect transmural flow in the cord tissue [25] and possible differences in solute concentration between the umbilical arterial blood and the amniotic fluid. The governing equations (see Supplement S1 for more details) in dimensionless variables describing solute transport in the umbilical tissue are given by
@@
$\nabla^{2}C$ $=\alpha\left(C+1/\beta\right)\quad\text{in }\;\mathcal{I}_{\text{t}}\,$, (1a)
$\eta\,C+(\mathbf{\hat{n}}\cdot\nabla C)$ $=0\quad\text{on }\;\mathcal{I}_{\text{c}}\,$, (1b)
$\mathbf{\hat{n}}\cdot\nabla C$ $=0\quad\text{on }\;\mathcal{I}_{\text{f,\,p}}\,$, (1c)
$C$ $=0\quad\text{on }\;\mathcal{I}_{\text{a${}_{i}$}}\;\text{for }\;i=1,2\,,\quad\text{and}\quad C=1\quad\text{on }\;\mathcal{I}_{\text{v}}\,.$ (1d)
Here $C(x,y,z)$ is the dimensionless solute concentration (or temperature) normalised by the vein-artery concentration difference $C_{v}-C_{a}$ and centred at the arterial concentration $C_{a}$, so that the dimensional concentration is $C_{a}+(C_{v}-C_{a})C$. In (1a), $\alpha$ is a dimensionless metabolic rate parameter (assumed zero for all the solutes but oxygen), and $\beta=(C_{v}-C_{a})/C_{a}$ is the relative solute concentration difference between the umbilical artery (UA) and the umbilical vein (UV) (see Supplement S1). Exchange at the interface $\mathcal{I}_{\text{c}}$ between the umbilical cord and the amniotic fluid is characterised by the amniotic exchange parameter $\eta$ in (1b); when $\eta=0$ no solute is lost through outer cord’s surface; $\eta\to\infty$ in (1b) enforces $C=0$ at the cord–amniotic interface. We impose no-flux conditions in (1c) at the tissue cross-sections bounding the fetal and placental ends of the umbilical cord respectively ($\mathcal{I}_{\text{f,\,p}}$); here $\mathbf{\hat{n}}$ denotes a unit normal pointing outside the domain occupied by the cord tissue $\mathcal{I}_{\text{t}}$. Finally, assuming zero axial concentration gradient along the cord, we set fixed solute concentrations on the surface of the umbilical arteries $\mathcal{I}_{\text{a${}_{1}$}}$, $\mathcal{I}_{\text{a${}_{2}$}}$ and the vein $\mathcal{I}_{\text{v}}$ in (1d), so that we can focus attention on intervascular exchange through cord tissue, accounting for its internal helical structure.
We reduce the three-dimensional problem (1) to two dimensions by exploiting the helical coordinate transformation $X=x\cos{\left(\Omega z\right)}+y\sin{\left(\Omega z\right)}$, $Y=-x\sin{\left(\Omega z\right)}+y\cos{\left(\Omega z\right)}$, $Z=z$, where $z$ measures the distance along the cord’s axis (see Supplement S1 for more details). This transformation allows us to study the three-dimensional effects of helicity in the cross-sectional plane of the cord by solving for $C(X,Y)$, the field that is invariant along the length of the cord relative to a rotating coordinate system. We assume that all vessels share the same pitch and that each vessel centreline maintains a fixed distance from the (straight) cord centreline. The dimensionless solute flux $N$ per unit length of the cord delivered from the UV to the two UAs and cord tissue, and the solute flux per unit length taken up by cord tissue are then given, respectively, by
$N(\Omega,\alpha,\beta,\eta)=\int_{\partial\mathcal{I}_{\text{v}}\,\cap\,\mathcal{P}}\mathbf{\hat{n}}\cdot\left(\nabla_{\perp}C+\Omega^{2}\mathbf{H}\right)\mathop{}\mathrm{d}s\quad\;\text{and}\;\quad N_{\text{u}}=\int_{\mathcal{I}_{\text{t}}\,\cap\,\mathcal{P}}\!\alpha\left(C+1/\beta\right)\!\mathop{}\mathrm{d}A\,$, (2)
where $\nabla_{\perp}\equiv[\partial_{X},\,\partial_{Y},\,0]$ is the 2D gradient operator, and $\mathbf{H}=(Y\,C_{X}-X\,C_{Y})\,[Y,\,-X,\,0]$ is a flux component arising due to cord helicity (see Supplement S2 for further details). The integrals in (2) are taken in a plane $\mathcal{P}$, in which $z=\text{constant}$. The solute flux $N(\Omega,\alpha,\beta,\eta)$ in transformed coordinates shows explicitly the rotational effects of helicity. The helicity factor $\Omega^{2}$ controls the strength of the rotational effects while $\mathbf{H}$ represents how the concentration gradients twist in space. This term shows that concentration not only diffuses radially but also experiences angular distortions relative to the vessels. We explore below the effects of helicity, vessel proximity and the configuration of the vessels within the cord on the solute fluxes. The ratio of a dimensional exchange flux between the umbilical vessels, based on $N$, and the advective flux along the umbilical cord defines a dimensionless Damköhler number
$\mathrm{Da}=\frac{D^{*}_{t}L^{*}\,N(\Omega,\alpha,\beta,\eta)}{B\,Q^{*}}\,$, (3)
where $D^{*}_{t}$ is the solute diffusivity in tissue, $L^{*}$ is the cord length, $Q^{*}$ is the total umbilical flow rate and $B$ is an advection-facilitation parameter that accounts for haemoglobin binding of certain solutes, such as oxygen; see Supplement S8 and Table S3 for more details and estimates of $\mathrm{Da}$ for oxygen and heat exchange.
@@ II.2 Image analysis of histology and ultrasound data
Histology and ultrasound images of human cord cross-sections were taken from published literature [7, 26, 27, 28, 18, 29, 30, 31]. The histology data (stained with hematoxylin and eosin; see Fig. 1b) corresponds to normal pregnancies, with gestational age varying from 37 to 42 weeks; ultrasound data included cords from the third trimester (see Supplementary Table S4 for more details on data sources).
@@ III.1 Cord coiling and vascular proximity promote solute and heat shunting
Figure 2: The impact of cord helicity and vascular proximity on solute exchange. (a) Relative exchange flux $N_{\mathrm{rel}}$ per unit length (4) of the umbilical vein vs. cord helicity $\Omega$ for different UA-UV separation distances $d=0.3$ and $d=0.04$, with $\alpha=0,\,\eta=0$. For weakly coiled cords, the solute (heat) exchange amplification scales as $\Omega^{2}$ (relative to the uncoiled vessels of the same length). Concentration fields for four different vessel configurations and three-dimensional diagrams show a normocoiled ($\Omega$=0.75) and hypercoiled ($\Omega$=3.77) cord, at parameters highlighted with squares on the graph. Cords with $\Omega<0.4$ are hypocoiled while cords with $\Omega>2$ are hypercoiled. The colour bar shows solute concentration in normalised units (venous concentration is unity and arterial concentration is zero). (b) Computed exchange flux per unit length of the vein NN for small separation $d$. In the limit of small $d$, the leading-order flux is $\mathcal{O}(d^{-1/2})$. Here $\alpha=0,\,\eta=0$. (c) Exchange flux NN and relative exchange flux $N_{\mathrm{rel}}$ for different values of the exchange parameter $\eta$ in the cord boundary condition (1b) for $d=0.33$. Other geometric parameters in (a-c) were fixed: $R_{v}=0.25$, $r_{v}=0.25$, $r_{a}=0.12$, $\theta_{1}=0.8\pi$, $\theta_{2}=1.2\pi$ (see Supplement for more details). The triangles in (a) and (b) show the relative fold-change in the variables.
@@
$N_{rel}=\frac{N(\Omega,\alpha,\beta,\eta)-N(0,\alpha,\beta,\eta)}{N(0,\alpha,\beta,\eta)},$ (4)
assuming initially that $\alpha=\eta=0$. Small $\Omega$ corresponds to weakly coiled cords, while a large $\Omega$ corresponds to a tightly coiled cord. This is shown in the 3D representations of one complete coil for $\Omega=0.75$ and $\Omega=3.77$ (Fig. 2a, insets). For large $\Omega$, $N_{rel}$ exceeds unity, implying that the flux can be twice as large as the flux for straight vessels, because tightly coiled cords provide a larger internal surface area for exchange to take place. The concentration field for $\Omega=3.77$, $d=0.3$ (Fig. 2a, inset, top right) varies weakly around the cord perimeter, showing how coiling promotes direct solute exchange between vessels. In contrast, the flux for weakly coiled vessels is similar to the flux for straight vessels. Fig. 2(a) shows that the first correction to the straight vessel flux is of order $\Omega^{2}$. A weak helicity $\Omega\ll 1$ approximation for flux can therefore be obtained by perturbing the three-dimensional problem (1a) in powers of $\Omega^{2}$.
Since in the problem for $C(X,Y)$ the helicity term appears as $\Omega^{2}$, there is no evidence of chirality in the concentration fields in Fig. 2(a). In the Supplement (Fig. S8b,c), we confirm that two configurations with opposite chirality have the same cross-sectional concentration fields. The position of the arteries in the configuration considered in Fig. 2(a) is symmetrical relative to the $x$-axis. This leads to the symmetry C(X,Y)=C(X,Y)C(X,Y)=C(X,-Y). The Supplement (Fig. S8a) shows concentration fields with increasing helicity for asymmetric configurations. Breaking the symmetry introduces distortion in the concentration fields, but the trend in relative flux is the same as in Fig. 2(a).
Another geometric parameter that contributes to the exchange of solutes is the separation distance $d$ between vessels. The flux is larger for the umbilical vessels that are closer together because concentration gradients are steeper, as shown in Fig. 2(a), colour map inserts. Indeed, Fig. 2(b) illustrates how, as $d$ gets smaller, the flux in the gap between the vessels of $\mathcal{O}(d^{-1/2})$ dominates over the flux outside the gap (see Supplement S3 for more details), and the weak helicity correction of $\mathcal{O}(d^{1/2}\,\Omega^{2})$ provides a good approximation ($d=0.04$; Fig. 2a). On the other hand, for larger $d$, the effect of helicity becomes more dominant, as demonstrated by the larger excess flux $N_{rel}$ (4) relative to the flux in the uncoiled geometry.
@@
For the range of exchange fluxes $N$ predicted in Fig. 2, the characteristic values of the Damköhler number (3), which quantifies the extent of ‘shunting’ between the UV and UA, are $\mathrm{Da}_{\text{O}_{2}}\sim 10^{-5}$ and $\mathrm{Da}_{\text{heat}}\sim 10^{-1}$ for oxygen and heat, respectively (see Supplement S8 for more details). Thus, although no physiological shunting is expected for oxygen exchange, heat exchange could be sensitive to the cord configuration.